OpenStudy (anonymous):

The regular price of a watch is $32.25. During a sale, the watch was marked 16% off. What was the price of the watch during the sale? $5.16 $16.25 $21.16 $27.09

OpenStudy (bloomlocke367):

Well, since it is 16% off, you know the person buying the watch only payed 84% of the original price. You could set up a proportion to find out how much he payed, like such: \(\LARGE\frac{1}{32.25}=\frac{0.86}{?}\) I changed it to decimals by dividing the percents by 100. Now you just cross-multiply and divide.

OpenStudy (sbuck98):

you do 32.25*16%=5.16 32.25-5.16=27.09

OpenStudy (bloomlocke367):

@sbuck98 please refrain from giving direct answers as they are against the code of conduct! :)

OpenStudy (anonymous):

b?

OpenStudy (sbuck98):

D

OpenStudy (bloomlocke367):

I meant to say 0.84 :O my bad

OpenStudy (sbuck98):

I gave the direct answer after I showed him how to work it, don' t get butt hurt I answered before you it's just a medal.... @BloomLocke367

OpenStudy (bloomlocke367):

@geny55 how did you get b? (it is incorrect) but I want to know your thinking to better help you :)

OpenStudy (anonymous):

actually sorry i got d

OpenStudy (anonymous):

next!

OpenStudy (anonymous):

If a sprinkler waters 1 over 14 of a lawn in 1 over 2 hour, how much time will it take to water the entire lawn? 1 over 7 hour 4 over 7 hour 4 hours 7 hours

OpenStudy (anonymous):

d?

OpenStudy (sbuck98):

I think so

OpenStudy (anonymous):

because 14* 0.5 = 7

OpenStudy (anonymous):

5.16$